eviant:

Solve the equation by factoring. c2 – 8c = 0

eviant:

@dude

dude:

\(c^2-8c=0\) What is a common factor in both of these terms?

eviant:

c

dude:

Right so we can do \(c(c-8)=0\)

dude:

Now you set the factors equal to 0 \(c=0\) and \(c-8=0\) Those are your solutions